OpenStudy (anonymous):

Write the slope -intercept from of the equation .. Through (2,-3),slope=1/2

OpenStudy (texaschic101):

y = mx + b slope(m) = 1/2 (2,-3) x = 2 and y = -3 we have everything given to us except for b, so we have to find b y = mx + b -3 = 1/2(2) + b -3 = 1 + b -3 - 1 = b -4 = b therefore, your equation is : y = 1/2x - 4